OpenStudy (anonymous):

Please help: Write in polar form 1-sqrt(3)*i

OpenStudy (anonymous):

\[1-\sqrt{3}i\]

OpenStudy (mathmath333):

\(\large \tt \begin{align} \color{black}{z=a+bi\\~\\ r=|z|=\sqrt{a^2+b^2}\\~\\ \alpha=tan^{-1}(\dfrac{b}{a})\\~\\ trig form \\~\\ z=r(cos\alpha+isin\alpha)\\~\\ } \end{align}\)

OpenStudy (anonymous):

But is there a way to simplify \[z=2(\cos( \frac{ \pi}{ 3 })+isin(\frac{ \pi }{ 3}))\]

OpenStudy (jdoe0001):

nope, that's the polar form

OpenStudy (anonymous):

Is this correct when the argument is between o and 2\[\pi \]?

OpenStudy (mathmath333):

there is mistake in ur polar form i think

OpenStudy (anonymous):

which part is the mistake?

OpenStudy (mathmath333):

the angle

OpenStudy (mathmath333):

\(\alpha\)

OpenStudy (anonymous):

\[\tan \prime(\sqrt{3})\]

OpenStudy (mathmath333):

\(tan^{-1}(-\sqrt{3})\)

OpenStudy (anonymous):

Oh it was negative...

OpenStudy (anonymous):

okay thank you!

OpenStudy (mathmath333):

so its \(\Large 2[cos(300^{\circ})+isin(300^{\circ})]\)

OpenStudy (mathmath333):

\(\Large 2[cos(\dfrac{5\pi}{3})+isin(\dfrac{5\pi}{3})]\)